OpenStudy (anonymous):

r/2 - 6 =14 answers 1.r=40 2.r=34 3.r=20 4.r=16

OpenStudy (anonymous):

It's R=40!

OpenStudy (anonymous):

1.

OpenStudy (harsha19111999):

Option 1

OpenStudy (solomonzelman):

`1) Add 6 to both sides ` `2) multiply both sides times 2. `

OpenStudy (anonymous):

TYHank guys

OpenStudy (anonymous):

Or you can reverse the equation and do (14+6)*2 =3

OpenStudy (anonymous):

I have another

OpenStudy (anonymous):

Go for it

OpenStudy (anonymous):

3x - 3 = -18

OpenStudy (anonymous):

Also ignore the =3, that was supposed to be an emote, not part of the equation.

OpenStudy (anonymous):

-5!

OpenStudy (anonymous):

(-18+3) / 3 is -5, input -5 to that original equation and its correct.

OpenStudy (anonymous):

answers 1.x = -17 2.x = -5 3.x = -3 4.x= 5

OpenStudy (anonymous):

3x-3=-18 Add 3 to both sides 3x=-18+3 3x=-15 divide by 3 x=-5

OpenStudy (anonymous):

2. X = -5!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x = -5

OpenStudy (anonymous):

I have another

OpenStudy (anonymous):

No more than 12 people can enter the hunted house every 10 minus. which inequality represent the situation

OpenStudy (anonymous):

answers 1.a<12 2.a <_ 12 3.a >12 4.a >_ 12

OpenStudy (anonymous):

Assuming A is amount of people, 2. A is less than or Equal to 12.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so a<_ 12

OpenStudy (anonymous):

what grade lvl is this?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

d - 6 > -14

OpenStudy (anonymous):

sorry wrong one

OpenStudy (anonymous):

D = -7 or higher! no problem lol

OpenStudy (anonymous):

5w <_ 15

OpenStudy (anonymous):

If you can, close this question and post a new one. I'll help you like that, its just so i can get some more credit for helping c:

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

w = 3 or any number less than 3!